OpenStudy (anonymous):

determine the midpoint of the segment joining the given two points 1. (3a-1, a) and ( 5a+5, -a )

OpenStudy (anonymous):

@jdoe0001 what about this??

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 3a-1}}\quad ,&{\color{blue}{ a}})\quad &({\color{red}{ 5a+5}}\quad ,&{\color{blue}{ -1}}) \end{array}\quad \left(\cfrac{{\color{red}{ (5a-5)}} + {\color{red}{(3a-1)}}}{2} , \cfrac{{\color{blue}{ -a}} + {\color{blue}{ a}}}{2} \right)\)

OpenStudy (jdoe0001):

hmmm I have an -1 there =) anyow.. just a typo

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 3a-1}}\quad ,&{\color{blue}{ a}})\quad &({\color{red}{ 5a+5}}\quad ,&{\color{blue}{ -a}}) \end{array}\qquad \left(\cfrac{{\color{red}{ (5a-5)}} + {\color{red}{(3a-1)}}}{2} , \cfrac{{\color{blue}{ -a}} + {\color{blue}{ a}}}{2} \right)\)

OpenStudy (anonymous):

,, would the answer be (8a-4)/2 and 0?

OpenStudy (jdoe0001):

well... \(\bf \left(\cfrac{{\color{red}{ (5a-5)}} + {\color{red}{(3a-1)}}}{2}\quad ,\quad \cfrac{{\color{blue}{ -a}} + {\color{blue}{ a}}}{2} \right)\implies \left(\cfrac{8a-6}{2}\quad ,\quad 0\right)\)

OpenStudy (anonymous):

its 5a+5

OpenStudy (jdoe0001):

ohh shoot.. it's =(

OpenStudy (jdoe0001):

then yes :) is \(\bf \left(\cfrac{{\color{red}{ (5a+5)}} + {\color{red}{(3a-1)}}}{2}\quad ,\quad \cfrac{{\color{blue}{ -a}} + {\color{blue}{ a}}}{2} \right)\implies \left(\cfrac{8a-4}{2}\quad ,\quad 0\right)\)

OpenStudy (jdoe0001):

wait ah sec dohh darn

OpenStudy (anonymous):

thanks , still have questions , :3

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{{\color{red}{ (5a+5)}} + {\color{red}{(3a-1)}}}{2}\quad ,\quad \cfrac{{\color{blue}{ -a}} + {\color{blue}{ a}}}{2} \right)\implies \left(\cfrac{8a+4}{2}\quad ,\quad 0\right)\) +5-1 = +4